# Fine-tuning an Large Language Model for Automating Computational Fluid Dynamics Simulations

### v1

@@ 3 Results
We evaluate our fine-tuned LLM against open-source LLMs (Qwen2.5-72B-Instruct [9], DeepSeek-R1 [10], Llama3.3-70B-Instruct [11]) and the RAG-based MetaOpenFOAM [16] (using GPT-4o [8]). To isolate the impact of base LLM performance, we integrated the open-source LLMs into our framework through component substitution, retaining the multi-agent architecture’s verification and error-correction modules. The LLM temperature was set to 0.7 for these open-source models. The sampling randomness parameter, known as “temperature” for LLMs, was set to 0.7 for these open-source models. For the MetaOpenFOAM comparison, we used equivalent mesh files and adopted its GPT-40 temperature setting of 0.01.
@@ 3.1 Overall Performance
Our benchmark findings demonstrate that domain-specific fine-tuning enables LLMs to effectively bridge natural language instructions and CFD simulation, aided by the multi-agent framework’s valuable error correction. Tracking correction iterations revealed most errors involved missing basic parameters, such as a pressure reference. While sometimes omitted by the fine-tuned LLM, the corrector agent typically resolved such issues efficiently. Analysis of failures showed most errors occurred in turbulence simulations, mainly due to divergence caused by unreasonable parameter values, such as dissipation rates in the $k$-$\omega$ model was set orders of magnitude too high. Future work will extend the dataset with more diverse examples to improve fine-tuning and reduce these failures.
@@ 4 Conclusion
We developed an approach to automate CFD simulations by fine-tuning a large language model on domain-specific data. We constructed NL2FOAM, a dataset of 28716 natural language-to-OpenFOAM configuration pairs with chain-of-thought reasoning annotations, and fine-tuned Qwen2.5-7B-Instruct using LoRA to reduce trainable parameters from 7.6B to 0.02B. This domain-specific adaptation enables the LLM to translate natural language descriptions into complete OpenFOAM configurations. This fine-tuned LLM serves as the core of a multi-agent system (including pre-checker, LLM generator, runner, and corrector agents) that orchestrates the simulation workflow, ensuring syntax compliance and numerical stability.
@@
This research introduces a new paradigm for engineering automation that bridges natural language interfaces with specialized numerical simulations. By allowing non-experts to configure CFD simulations through natural language descriptions, our approach democratizes access to simulation capabilities while maintaining high accuracy. The multi-agent framework further demonstrates how domain-specific LLMs can be integrated with verification and correction mechanisms to achieve reliable automation of technically complex workflows requiring both linguistic understanding and numerical precision.
While our method performs well for incompressible benchmarks, limitations remain for more complex simulations. Future work will focus on expanding the NL2FOAM dataset to include more complex transport phenomena, e.g., multiphase flows, compressible flows, and heat transfer problems. We also plan to explore fine-tuning larger base models while maintaining computational efficiency, to enhance robustness and handle a wider variety of cases. These advancements will further broaden the approach’s applicability for engineering challenges.

### v2

@@ 3 Results
We evaluated our fine-tuned LLM against open-source LLMs (Qwen2.5-72B-Instruct [9], DeepSeek-R1 [10], and Llama3.3-70B-Instruct [11]) and the RAG-based MetaOpenFOAM [16] (using GPT-4o [8]). To isolate the impact of base LLM performance, we integrated the open-source LLMs into our framework through component substitution, retaining the multi-agent architecture’s verification and error-correction modules. The sampling randomness parameter, known as “temperature” for LLMs, was set to 0.7 for these open-source models. For the MetaOpenFOAM comparison, we used the same mesh files and adopted its GPT-4o temperature setting of 0.01.
@@ 3.1 Overall Performance
Benchmark demonstrates that larger general-purpose models, lacking specialized CFD knowledge, frequently generate incomplete or inadequate configurations. Our analysis showed they often omit necessary files or parameters (e.g., solver-specific settings) and produce parameter values lacking the diversity and flexibility needed for varied CFD scenarios. This inflexibility, particularly with different turbulence models or flow conditions, results in higher simulation failure or divergence rates.
In contrast, domain-specific fine-tuning enables the model to effectively bridge natural language instructions and the OpenFOAM setup. This results in more relevant and robust configurations tailored to CFD tasks, confirmed by the lower iteration metric. Tracking correction iterations revealed that most errors involved missing basic parameters, such as a pressure reference, which the corrector agent resolved efficiently. The most persistent challenges occur in turbulence simulations, mainly due to divergence caused by unreasonable parameter values, such as dissipation rates in the $k$-$\omega$ model was set orders of magnitude too high. Future work will extend the dataset with more diverse examples to improve fine-tuning and reduce these failures.
@@ 4 Conclusion
We developed an approach to automate CFD simulations by fine-tuning an LLM on domain-specific data. We constructed NL2FOAM, a dataset of 28716 natural language-to-OpenFOAM configuration pairs with chain-of-thought reasoning annotations, and fine-tuned Qwen2.5-7B-Instruct using LoRA to reduce trainable parameters from 7.6B to 0.02B. This domain-specific adaptation enables the LLM to translate natural language descriptions into complete OpenFOAM configurations. This fine-tuned LLM serves as the core of a multi-agent system (including pre-checker, LLM generator, runner, and corrector agents) that orchestrates the simulation workflow, ensuring syntax compliance and numerical stability.
@@
This research introduces a new paradigm for engineering automation that bridges natural language interfaces with specialized numerical simulations. By allowing non-experts to configure CFD simulations through natural language descriptions, our approach democratizes access to simulation capabilities while maintaining high accuracy. The multi-agent system further demonstrates how domain-specific LLMs can be integrated with verification and correction mechanisms to achieve reliable automation of technically complex workflows requiring both linguistic understanding and numerical precision.
While our method performs well on incompressible benchmarks, limitations remain for more complex simulations. Future work will focus on expanding the NL2FOAM dataset to include more complex transport phenomena, e.g., multiphase flows, compressible flows, and heat transfer problems, which involve additional configuration files and different solver setups. Computational efficiency and accuracy should be assessed when scaling to industrial-level problems for practical application. Furthermore, we plan to explore fine-tuning larger base models to enhance robustness and handle a wider variety of cases. These advancements will further broaden the approach’s applicability for engineering challenges.
@@ Acknowledgments
Appendix A Fine-tuning details
We applied LoRA with a rank $r=8$, reducing trainable parameters from 7.6B to 0.02B. Fine-tuning utilized Llamma-Factory [31], a unified framework enabling efficient and flexible LLM adaptation. We trained the model using four NVIDIA GeForce RTX 4090 GPUs, employing AdamW [32] optimization. The baseline learning rate was $5\times 10^{-5}$, with linear warmup over the first 10% of steps to mitigate initial instability. A total batch size of 16 balanced GPU memory constraints and training efficiency.
Training proceeded for four epochs, with checkpoints saved after each. This duration was empirically chosen based on validation loss plateaus to balance convergence and cost. As shown in Fig. 1, performance pass@1 peaked at 82.6% after two epoches, then declined slightly while remained above 75%, suggesting potential overfitting. Consequently, we selected the epoch two checkpoint for experiments.
Figure 1: Evolution of the performance pass@1 across epochs on our benchmark. The model reached a peak pass@1 of 82.6% at epoch two.